# An Improved Privacy and Utility Analysis of Differentially Private SGD with Bounded Domain and Smooth Losses

### v2

@@ Appendix C Proofs for Utility Analysis
where the last step is by Lemma A.13. Substituting Equation 54 and Equation 55 into Equation 50, we have
$\mathbb{E}_{\boldsymbol{\zeta}_{t},\xi}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ $\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{22\eta c}{25\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}+\frac{\eta c}{\sqrt{2L}}\frac{2\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}}{5}+d\eta^{2}\sigma_{\text{DP}}^{2}$ (56)
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]\leq\mathcal{O}\left(\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{d\eta\sigma_{\text{DP}}^{2}\sqrt{L}}{C}\right).$ (57)
@@
By using Lemma A.14, we immediate obtain $\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\leq\sqrt{\frac{1}{2\mu}}\|\nabla l(\boldsymbol{\theta}_{t})\|\leq\mathcal{O}\left(\frac{L^{3/4}}{\mu^{5/4}}\sigma_{\text{SGD}}\right)$.
@@
$\mathbb{E}_{\boldsymbol{\zeta}_{t}}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ $\leq\mathbb{E}_{\zeta_{t}}\left[\|\boldsymbol{\theta}_{t}-\eta g(\boldsymbol{\theta}_{t})+\boldsymbol{\zeta}_{t}-\boldsymbol{\theta}^{*}\|^{2}\right]$ (58)
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})+\eta^{2}\|g(\boldsymbol{\theta}_{t})\|^{2}+d\eta^{2}\sigma_{\text{DP}}^{2},$
@@
$\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$ $=\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma\nabla l(\boldsymbol{\theta}_{t})+\gamma\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$ (59)
$\leq-2\eta\gamma[\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})-\mathbb{E}_{\xi}\left[2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$
$\leq-\frac{\eta C}{\|\nabla l(\boldsymbol{\theta}_{t})\|}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]+\mathbb{E}_{\xi}\left[2\eta\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|\right]$
@@
where the third step is by the convexity and Cauchy-Schwarz inequality, and the fourth step is by Lemma A.13 and Lemma A.14. Combined with Equation 55, we have
$\mathbb{E}_{\boldsymbol{\zeta}_{t},\xi}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ $\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{3\eta C}{5\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}+\frac{\eta C}{\sqrt{2L}}\frac{2\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}}{5}+\eta^{2}\sigma_{\text{DP}}^{2}$ (60)
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]\leq\mathcal{O}\left(\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{d\eta\sigma_{\text{DP}}^{2}\sqrt{L}}{C}\right).$ (61)
@@
Let $\kappa=\mathbbm{1}\{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})\|>C\}$ and use Lemma A.1, yielding
$\operatorname{Pr}\left[\kappa=1\right]\leq\operatorname{Pr}\left[\{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|^{2}>\frac{C^{2}}{4}\}\right]\leq\frac{4\sigma_{\text{SGD}}^{2}}{C}.$ (63)
@@
$\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$ $=-2\eta\left[\mathbb{E}_{\xi}[g_{\xi}(\boldsymbol{\theta}_{t})]-\nabla l(\boldsymbol{\theta}_{t})+\nabla l(\boldsymbol{\theta}_{t})\right]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$ (64)
$\leq 2\eta\cdot\|\mathbb{E}_{\xi}[g_{\xi}(\boldsymbol{\theta}_{t})]-\nabla l(\boldsymbol{\theta}_{t})\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|-2\eta\nabla l(\boldsymbol{\theta}_{t})^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$\leq 2\eta\cdot\|\mathbb{E}_{\xi}\left[g_{\xi}(\boldsymbol{\theta}_{t})-\nabla l_{\xi}(\boldsymbol{\theta}_{t})|\kappa=1\right]\cdot\operatorname{Pr}(\kappa=1)\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$\leq\frac{8\eta\sigma_{\text{SGD}}^{2}}{C^{2}}\|\mathbb{E}_{\xi}[(1-\frac{C}{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})\|})\nabla l_{\xi}(\boldsymbol{\theta}_{t})]\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$\leq\frac{8\eta\sigma_{\text{SGD}}^{2}}{C^{2}}\mathbb{E}_{\xi}[||\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})+\nabla l(\boldsymbol{\theta}_{t})||]\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$\leq\frac{4\eta\sigma_{\text{SGD}}^{3}}{\mu C}+\frac{16\eta\sigma_{\text{SGD}}^{2}}{C^{2}}\sqrt{\frac{L}{\mu}}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
@@
where the second step is by the convexity and Cauchy-Schwarz inequality, the third step is by the law of total expectation, the fifth step is by Jensen’s inequality, and the third to last step is by Lemma A.13 and Lemma A.14.
@@
$\mathbb{E}\left[\eta^{2}\|g(\boldsymbol{\theta}_{t})\|^{2}\right]$ $\leq\frac{\eta^{2}}{b^{2}}\sum_{\xi\in\mathcal{B}_{t}}\mathbb{E}_{\xi}\left[\|g_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})+l(\boldsymbol{\theta}_{t})\|^{2}\right]$ (65)
@@
Substituting Equation 64 and Equation 65 into Equation 58, we have
$\mathbb{E}_{\boldsymbol{\zeta}_{t},\xi}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ $\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}+\frac{4\eta\sigma_{\text{SGD}}^{3}}{\mu C}-\frac{46\eta}{25}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]+\frac{2\eta^{2}\sigma_{\text{SGD}}^{2}}{b}+\frac{9\eta}{5}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]+d\eta^{2}\sigma_{\text{DP}}^{2}$ (66)
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}+\frac{4\eta\sigma_{\text{SGD}}^{3}}{\mu C}-\frac{\eta}{25}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]+\frac{2\eta^{2}\sigma_{\text{SGD}}^{2}}{b}+d\eta^{2}\sigma_{\text{DP}}^{2}.$
@@
$\frac{1}{25(T+1)}\sum_{t=0}^{T}\mathbb{E}\left[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})\right]\leq\frac{D^{2}}{\eta(T+1)}+\frac{4\sigma_{\text{SGD}}^{3}}{\mu C}+\frac{2\eta\sigma_{\text{SGD}}^{2}}{b}+d\eta\sigma_{\text{DP}}^{2}.$ (67)
Exploiting Lemma A.4 and Lemma A.5, one may get
$\frac{1}{25(T+1)}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]$ $\leq\sqrt{\frac{D^{2}}{\eta(T+1)}+\frac{4\sigma_{\text{SGD}}^{3}}{\mu c}+\frac{2\eta\sigma_{\text{SGD}}^{2}}{b}+d\eta\sigma_{\text{DP}}^{2}}$ (68)
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]\leq\mathcal{O}\left(\frac{D}{\sqrt{\eta T}}+\sqrt{\frac{\sigma_{\text{SGD}}^{3}}{\mu C}}+\frac{\sqrt{\eta}\sigma_{\text{SGD}}}{\sqrt{b}}+\sqrt{d\eta}\sigma_{\text{DP}}\right).$ (69)

### v3

@@ Appendix C Proofs for Utility Analysis
where the last step is by Lemma A.14. Substituting (62) and (63) into (58), we have
$\mathbb{E}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ (64)
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{22\eta c}{25\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
$+\frac{\eta c}{\sqrt{2L}}\frac{2\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}}{5}+d\eta^{2}\sigma_{\text{DP}}^{2}$
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]$ (65)
$\leq\mathcal{O}\left(\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{d\eta\sigma_{\text{DP}}^{2}\sqrt{L}}{C}\right).$
@@
By using Lemma A.15, we immediate obtain
$\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\leq\sqrt{\frac{1}{2\mu}}\|\nabla l(\boldsymbol{\theta}_{t})\|\leq\mathcal{O}\left(\frac{L^{3/4}}{\mu^{5/4}}\sigma_{\text{SGD}}\right).$ (66)
@@
$\mathbb{E}_{\boldsymbol{\zeta}_{t}}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ (67)
$\leq\mathbb{E}_{\zeta_{t}}\left[\|\boldsymbol{\theta}_{t}-\eta g(\boldsymbol{\theta}_{t})+\boldsymbol{\zeta}_{t}-\boldsymbol{\theta}^{*}\|^{2}\right]$
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$+\eta^{2}\|g(\boldsymbol{\theta}_{t})\|^{2}+d\eta^{2}\sigma_{\text{DP}}^{2},$
@@
$\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$ (68)
$=\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma\nabla l(\boldsymbol{\theta}_{t})+\gamma\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$
$\leq-2\eta\gamma[\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$-\mathbb{E}_{\xi}\left[2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$
$\leq-\frac{\eta C}{\|\nabla l(\boldsymbol{\theta}_{t})\|}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$+\mathbb{E}_{\xi}\left[2\eta\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|\right]$
@@
where the third step is by the convexity and Cauchy-Schwarz inequality, and the fourth step is by Lemma A.14 and A.15. Combined with (63), we have
$\mathbb{E}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ (69)
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{3\eta C}{5\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
$+\frac{\eta C}{\sqrt{2L}}\frac{2\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}}{5}+\eta^{2}\sigma_{\text{DP}}^{2}$
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]$ (70)
$\leq\mathcal{O}\left(\frac{\sqrt{L}D^{2}}{\eta CT}+\frac{d\eta\sigma_{\text{DP}}^{2}\sqrt{L}}{C}\right).$
@@
Let $\kappa=\boldsymbol{1}\{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})\|>C\}$ and use Lemma A.2, yielding
$\operatorname{Pr}\left[\kappa=1\right]$ $\leq\operatorname{Pr}\left[\{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|^{2}>\frac{C^{2}}{4}\}\right]$ (72)
$\leq\frac{4\sigma_{\text{SGD}}^{2}}{C}.$
@@
$\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$ (73)
$=-2\eta\left[\mathbb{E}_{\xi}[g_{\xi}(\boldsymbol{\theta}_{t})]-\nabla l(\boldsymbol{\theta}_{t})+\nabla l(\boldsymbol{\theta}_{t})\right]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$\leq 2\eta\cdot\|\mathbb{E}_{\xi}[g_{\xi}(\boldsymbol{\theta}_{t})]-\nabla l(\boldsymbol{\theta}_{t})\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|$
$-2\eta\nabla l(\boldsymbol{\theta}_{t})^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$\leq 2\eta\cdot\|\mathbb{E}_{\xi}\left[g_{\xi}(\boldsymbol{\theta}_{t})-\nabla l_{\xi}(\boldsymbol{\theta}_{t})|\kappa=1\right]\cdot\operatorname{Pr}(\kappa=1)\|$
$\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$\leq\frac{8\eta\sigma_{\text{SGD}}^{2}}{C^{2}}\|\mathbb{E}_{\xi}[(1-\frac{C}{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})\|})\nabla l_{\xi}(\boldsymbol{\theta}_{t})]\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|$
$-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
[…]
$-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$\leq\frac{4\eta\sigma_{\text{SGD}}^{3}}{\mu C}+\frac{16\eta\sigma_{\text{SGD}}^{2}}{C^{2}}\sqrt{\frac{L}{\mu}}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$-2\eta[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
@@
where the second step is by the convexity and Cauchy-Schwarz inequality, the third step is by the law of total expectation, the fifth step is by Jensen’s inequality, and the third to last step is by Lemma A.14 and A.15.
@@
$\mathbb{E}\left[\eta^{2}\|g(\boldsymbol{\theta}_{t})\|^{2}\right]$ (74)
$\leq\frac{\eta^{2}}{b^{2}}\sum_{\xi\in\mathcal{B}_{t}}\mathbb{E}_{\xi}\left[\|g_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})+l(\boldsymbol{\theta}_{t})\|^{2}\right]$
@@
Substituting (73) and (74) into (LABEL:eq:converge-4), we have
$\mathbb{E}_{\boldsymbol{\zeta}_{t},\xi}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ (75)
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}+\frac{4\eta\sigma_{\text{SGD}}^{3}}{\mu C}-\frac{46\eta}{25}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$+\frac{2\eta^{2}\sigma_{\text{SGD}}^{2}}{b}+\frac{9\eta}{5}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]+d\eta^{2}\sigma_{\text{DP}}^{2}$
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}+\frac{4\eta\sigma_{\text{SGD}}^{3}}{\mu C}-\frac{\eta}{25}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$+\frac{2\eta^{2}\sigma_{\text{SGD}}^{2}}{b}+d\eta^{2}\sigma_{\text{DP}}^{2}.$
@@
$\frac{1}{25(T+1)}\sum_{t=0}^{T}\mathbb{E}\left[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})\right]$ (76)
$\leq\frac{D^{2}}{\eta(T+1)}+\frac{4\sigma_{\text{SGD}}^{3}}{\mu C}+\frac{2\eta\sigma_{\text{SGD}}^{2}}{b}+d\eta\sigma_{\text{DP}}^{2}.$
Exploiting Lemma A.5 and A.6, one may get
$\frac{1}{25(T+1)}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]$ (77)
$\leq\sqrt{\frac{D^{2}}{\eta(T+1)}+\frac{4\sigma_{\text{SGD}}^{3}}{\mu c}+\frac{2\eta\sigma_{\text{SGD}}^{2}}{b}+d\eta\sigma_{\text{DP}}^{2}}$
@@
$\frac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\left[\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\right]$ (78)
$\leq\mathcal{O}\left(\frac{D}{\sqrt{\eta T}}+\sqrt{\frac{\sigma_{\text{SGD}}^{3}}{\mu C}}+\frac{\sqrt{\eta}\sigma_{\text{SGD}}}{\sqrt{b}}+\sqrt{d\eta}\sigma_{\text{DP}}\right).$